A steel bar 1.5m long is used to lift a stack of plates. If the fulcrum is 5 cm from the end of the bar and a force of 500 N is used, calculate A) the MA, B)the resistance force.
1
Expert's answer
2019-03-22T10:40:56-0400

Given:

"l=1.5~\text{m};~a =0.05~\text{m};~F=500~\text{N}. \\"

The mechanical advantage:

"MA = \frac{l-a}{a} = \frac{F_{resist}}{F}; \\" thus:

A)

"MA = \frac{1.5-0.05}{0.05} = 29.\\"

B)

"F_{resist} = F\frac{l-a}{a} = 500~\text{N}*\frac{1.5-0.05}{0.05} = 14500~\text{N}. \\"